# Eigenstate Thermalization Hypothesis A Short Review

### v1

@@ A Short Review
Understanding how an isolated quantum system evolves toward a thermal state from an initial state far from equilibrium such as one prepared by a global quantum quench has attracted significant interest in recent years. This phenomenon can be elucidated through the Eigenstate Thermalization Hypothesis (ETH), which has had a profound impact across various fields, from high-energy physics to condensed matter physics. The purpose of this review article is to present the fundamental concepts of quantum equilibrium and the ETH to a broad audience within the physics community, particularly for those in high-energy physics who seek a comprehensive understanding of these important topics.
Based on the review talk presented at the ”Workshop on Dynamics and Scrambling of Quantum Information”, December 2024, IPM.
@@ 5 The role of initial state: Weak and Strong thermalization
Using the analytic expression for the expectation value of energy, we have plotted the energy density, $\frac{E}{N}$, in Figure 5 for $N=100$ for a model characterized by the parameters $J_{z}=-1,h_{x}=1.05,h_{z}=-0.5$, with all other parameters set to zero, resulting in a non-integrable system. To emphasize the regions where the energy density approaches zero, we have depicted its absolute value. In this figure, the dark regions indicate initial states that are likely to exhibit strong thermalization. As one moves toward the lighter regions, the degree of thermalization decreases. In particular, the weakest thermalization occurs for the state $|\frac{\pi}{6},\pi\rangle$. It is worth mentioning that there are several other metrics available to assess weak or strong thermalization (see for example [22, 23, 24, 25, 26, 27]).
@@ Appendix B Mathematica codes
identity = IdentityMatrix[2]; sigamAtN[mat_,n_,L_] := Module[{list}, list = Table[identity, {L}]; list[[n]] = mat; Fold[KroneckerProduct, First[list], Rest[list]]];
@@
H[L_, g_, h_] := -g Sum[sigmax[i, L], {i, 1, L}] - h Sum[sigmaz[i, L], {i, 1, L}] - Sum[sigmaz[i, L] sigmaz[i + 1, L], {i, 1, L - 1}]
@@
initialState[theta_, phi_, L_] := Flatten[tensorProductState[singleState[theta, phi], L]]
@@
OinEVecH = ParallelTable[ ConjugateTranspose[EVectorH[[m]]].Sx[L].EVectorH[[n]],{m, 1, Length[EVectorH]},{n, 1, Length[EVecH]}];
@@
EValueH = Eigenvalues[N[H, 10]]; EVectorH= Eigenvectors[N[H, 10]];
@@
OinEVecH = ParallelTable[ N[Dot[ConjugateTranspose[EVectorH[[n]]].Sx.EVectorH[[m]]]], {m, 1, Length[EVectorH]}, {n, 1, Length[EVectorH]}];
@@
Do[Do[ listEV[theta, phi] = {}; listc = {}; For[m = 1, m <= Length[EValueH], m += 1, AppendTo[listc, {m, N[Dot[ConjugateTranspose[EVectorH[[m]]] . initialState[theta, phi, L]]]}]]; EVO[t_] := ParallelSum[ E^(I (EValueH[[m]] - EValueH[[n]]) t)* Conjugate[listc[[n, 2]]]* listc[[m, 2]] *OinEVecH[[m, n]], {n, 1, Length[EValueH]}, {m, 1, Length[EValueH]}] ; For[t = 0, t <= 30, t += 2/10, AppendTo[listEV[theta, phi], {t, EVO[t]}]], {phi, {Pi/2}}], {theta, {Pi/2}}]
@@
ListLinePlot[{listEV[Pi/2, Pi/2], {{0, 0}, {30, 0}}}, FrameTicks -> {{Automatic, None}}, InterpolationOrder -> 2, PlotStyle -> {{Blue}, {Orange, Dashed}}, PlotRange -> All, Frame -> True, FrameLabel -> {”t”, ”< \!\(\*SubscriptBox[\(S\), \(x\)]\) >”}, LabelStyle -> Directive[FontSize -> 10], RotateLabel -> True, FrameStyle -> Directive[Black, Bold] ]
@@
EValueH = Eigenvalues[N[H, 10]]; EValueHbeta =-beta*Eigenvalues[N[H, 10]];
@@
Trcanonical= ParallelSum[ EValueH[[i]] E^(EValueHbeta[[i]]), {i, 1, Length[EValueH]}]/ ParallelSum[E^(EValueHbeta), {i, 1, Length[EValueH]}]
@@
AbsInversetemp = Abs[NSolve[Trcanonical == N[Tr[Dot[H, Dot[initialState[theta, phi, L] .ConjugateTranspose[initialState[theta, phi, L]]]]]], beta, Reals][[1, 1, 2]]]
@@
SPINQBasis = Permutations[{1, 1, 1, , 1, 1, 1, 0, 0, 0, …,0, 0, 0}, {L}];
@@
For[i = 1, i <= 2^L, i += 1, If[ 0.8<Sum[1/4(sortedEigensystem[[2, i, j]]+sortedEigensystem[[2, i, pair[j]]])^2, {j, 1, 2^L}]<1.2,parity[i] = 1, parity[i] = -1]]
@@
listE = {}; listO = {}; For[i = 1, i <= 2^L, i += 1, If[parity[i] == 1, AppendTo[listE, sortedEigensystem[[1, i]]], AppendTo[listO, sortedEigensystem[[1, i]]]]]
@@
EigenValueEven={} For[m = 1, m <= Length[leveleven], m += 1, AppendTo[EigenValueEven, (leveleven[[m]])/Mean[leveleven]]];
@@
[…]
EigenValueOdd={} For[m = 1, m <= Length[levelodd], m += 1, AppendTo[EigenValueOdd, (levelodd[[m]])/Mean[levelodd]] ];

### v2

@@ A Short Review
Understanding how an isolated quantum system evolves toward a thermal state from an initial state far from equilibrium—such as one prepared by a global quantum quench—has attracted significant interest in recent years. This phenomenon can be elucidated through the Eigenstate Thermalization Hypothesis (ETH), which has had a profound impact across various fields, from high-energy physics to condensed matter physics. The purpose of this review article is to present the fundamental concepts of quantum equilibrium and the ETH to a broad audience within the physics community, particularly for those in high-energy physics who seek a comprehensive understanding of these important topics.
Based on the review talk presented at the ”Workshop on Dynamics and Scrambling of Quantum Information,” December 2024, IPM.
@@ 5 The role of initial state: Weak and Strong thermalization
Using the analytic expression for the expectation value of energy, we have plotted the energy density, $\frac{E}{N}$, in Figure 5 for $N=100$ for a model characterized by the parameters $J_{z}=-1,h_{x}=1.05,h_{z}=-0.5$, with all other parameters set to zero, resulting in a non-integrable system. To emphasize the regions where the energy density approaches zero, we have depicted its absolute value. In this figure, the dark regions indicate initial states that are likely to exhibit strong thermalization. As one moves toward the lighter regions, the degree of thermalization decreases. In particular, the weakest thermalization occurs for the state $|\frac{\pi}{6},\pi\rangle$. It is worth mentioning that there are several other metrics available to assess weak or strong thermalization (see for example [22, 23, 24, 25, 26, 27, 28]).
@@ Appendix B Mathematica codes
identity = IdentityMatrix[2]; sigmaAtN[mat_, n_, L_] := Module[{list}, list = Table[identity, {L}]; list[[n]] = mat; Fold[KroneckerProduct, First[list], Rest[list]]];
@@
H[L_, g_, h_] := -g Sum[sigmax[i, L], {i, 1, L}] - h Sum[sigmaz[i, L], {i, 1, L}] - Sum[sigmaz[i, L] . sigmaz[i + 1, L], {i, 1, L - 1}];
@@
initialState[theta_, phi_, L_] := tensorProductState[singleState[theta, phi], L]
@@
OinEVecH = ParallelTable[ ConjugateTranspose[EVectorH[[m]]] . Sx[L] . EVectorH[[n]], {m, 1, Length[EVectorH]}, {n, 1, Length[EVectorH]}];
@@
EValueH = Eigenvalues[N[H[L, g, h], 10]]; EVectorH = Eigenvectors[N[H[L, g, h], 10]];
@@
OinEVecH = ParallelTable[ N[Dot[ConjugateTranspose[EVectorH[[n]]] . Sx[L] . EVectorH[[m]]]], {m, 1, Length[EVectorH]}, {n, 1, Length[EVectorH]}];
@@
Do[Do[listEV[theta, phi] = {}; listc = {}; For[m = 1, m <= Length[EValueH], m += 1, AppendTo[ listc, {m, N[Dot[ConjugateTranspose[EVectorH[[m]]] . Flatten[initialState[theta, phi, L]]]]}]]; EVO[t_] := ParallelSum[ E^(I (EValueH[[m]] - EValueH[[n]]) t)*Conjugate[listc[[n, 2]]]* listc[[m, 2]]*OinEVecH[[m, n]], {n, 1, Length[EValueH]}, {m, 1, Length[EValueH]}]; For[t = 0, t <= 30, t += 2/10, AppendTo[ listEV[theta, phi], {t, EVO[t]}]], {phi, {Pi/2}}], {theta, {Pi/ 2}}];
@@
ListLinePlot[{listEV[Pi/2, Pi/2], {{0, 0}, {30, 0}}}, FrameTicks -> {{Automatic, None}}, InterpolationOrder -> 1, PlotStyle -> {{Blue}, {Orange, Dashed}}, PlotRange -> All, Frame -> True, FrameLabel -> {”t”, ”< \!\(\*SubscriptBox[\(S\), \(x\)]\) >”}, LabelStyle -> Directive[FontSize -> 10], RotateLabel -> True, FrameStyle -> Directive[Black, Bold]]
@@
EValueH = Eigenvalues[N[H[L, g, h]]]; EValueHbeta = -beta*Eigenvalues[N[H[L, g, h]]];
@@
Trcanonical = ParallelSum[ EValueH[[i]] E^EValueHbeta[[i]], {i, 1, Length[EValueH]}]/ ParallelSum[E^EValueHbeta[[i]], {i, 1, Length[EValueH]}];
@@
rho[theta_, phi_, L_] := Dot[initialState[theta, phi, L] . ConjugateTranspose[initialState[theta, phi, L]]]; AbsInversetemp = Abs[NSolve[ Trcanonical == N[Tr[Dot[H[L, g, h], rho[theta, phi, L]]]], beta, Reals][[1, 1, 2]]]
@@
eigensystem = Eigensystem[N[H[L, g, h]]]; sortedEigensystem = Transpose[SortBy[Transpose[eigensystem], First]]; SPINQBasis = Permutations[{1, 1, 1, , 1, 1, 1, 0, 0, 0, …,0, 0, 0}, {L}];
@@
For[i = 1, i <= 2^L, i += 1, If[0.8 < Sum[ 1/4 (sortedEigensystem[[2, i, j]] + sortedEigensystem[[2, i, pair[j]]])^2, {j, 1, 2^L}] < 1.2, parity[i] = 1, parity[i] = -1]];
@@
listE = {}; listO = {}; For[i = 1, i <= 2^L, i += 1, If[parity[i] == 1, AppendTo[listE, sortedEigensystem[[1, i]]], AppendTo[listO, sortedEigensystem[[1, i]]]]];
@@
EigenValueEven = {}; For[m = 1, m <= Length[leveleven], m += 1, AppendTo[EigenValueEven, (leveleven[[m]])/Mean[leveleven]]];
@@
levelodd = {}; Delta = Floor[Sqrt[Length[listE]]]; For[i = 10 Delta, i <= Length[listO] - 10 Delta, i += 1, AppendTo[levelodd, 1/(listO[[i + Delta]] - listO[[i - Delta]]) (listO[[i + 1]] - listO[[i]])]];
EigenValueOdd = {}; For[m = 1, m <= Length[levelodd], m += 1, AppendTo[EigenValueOdd, (levelodd[[m]])/Mean[levelodd]]];